A bullet travelling with a velocity of 16 m/s penetrates a tree trunk and comes to rest in 0.4 m. Find the time taken during the retardation.

Grade:11

3 Answers

Kevin Nash
9 years ago
Sol. u = 16 m/s (initial), v = 0, s = 0.4 m. Deceleration a = (v^2-u^2)/2s = -320 m/s2. Time = t = (V -u)/a = (0-16)/(-320 ) = 0.05 sec.
Sai Sriker Reddy Kuskuntla
6 years ago
u=16 m/s v=0S=0.4A=(v*v)-(u*u)/2A=0-256/0.8=-320t =(v-u)/a= -16/-320=0.05 secondTime taken during its motion in trunk=0.05 secondRetradation =320 m/s
Sakshi Kumari
5 years ago
Initial velocity= u= 16 m/s
distance covered = S=  0.4 m
Final velocity=v= 0 m/s 
the equation of motion;
v²=u²+2aS
0=16²+2 x a x 0.4 
a=-16²/0.8  
a=-320 m/s
As it carries a negative sign it represent retardation. 
Time taken during the retardation is given by the equation;
 v=u+at 0= 16-320 x t   
t=16/320 
t=1/20 
t=0.05 s